How to solve linear equations using any method.
FOr this one I tried using the substitution method, but i always get stuck.
2x-5y=3
x-2y=7 I know i have to get one variable by itself and i can pick which one. WHere I am getting stuck is once I get Y =8 for the first one, when I still have X to deal with, from there on, I do not know what to do.

Answer by macston(5194)

Without seeing what you tried, it is hard to say where you went wrong, but it is simplest to solve the second equation for x. After you have solved for 1 variable, plug it back in to solve for the other.
x-2y=7 Add 2y to each side.
x=2y+7
2x-5y=3 Substitute for x
2(2y+7)-5y=3
4y+14-5y=3
14-y=3 Add y to each side
14=y+3 Subtract 3 from each side
11=y ANSWER 1: y=11
x=2y+7=2(11)+7=29 ANSWER 2: x=29
CHECK:
2x-5y=3
2(29)-5(11)=3
58-55=3
3=3
x-2y=7
29-2(11)=7
29-22=7
7=7